OpenStudy (anonymous):

Simplify -20 - ( -5) ( 4)

OpenStudy (anonymous):

Remember order of operations here. What is -5 multiplied by +4?

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

What is -(-20) ?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

What is -20 + 20?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

\[-20-(-5)(4)=-20-(-20)=-20+20=0\]

OpenStudy (anonymous):

There you are.

OpenStudy (anonymous):

I assume the same answer for -20 - (-5) (-2 squared)